### v1

@@ 4.5 Symmetries with differential terms $Q_{(n)}$ (36) quadratic in spatial variables
The last class of symmetries which we are supposed to find are integrals of motion (7) generated by Killing tensors (39). The specificity of this class is that it includes integrals of motions with different transformation properties with respect to the discrete symmetry (25). Namely, the vector integrals of motion generated by $\mu_{(1)}^{ab}$ are invariant with respect to transformations (25) while the tensor integrals of motion generated by $\mu_{(2)}^{ab}$ change their sign under this transformation. It means that in fact we have two independent subclasses of integrals of motion since the linear combinations of integrals of motion with different transformations properties with respect to symmetry transformations are forbidden.
@@
Let us start with vector integrals of motion. Up to rotation transformations the only nonzero parameter in $\mu_{(1)}^{ab}$ (39) is $\lambda^{3}$, and the related equations (52) are reduced to the following form:
@@
Using relation (64) and the following condition
$x_{a}N_{a}=0$ (71)
(which is a consequence of (70)) equations (70) can be reduced to the following system including two equations only:
$\begin{split}&x_{1}^{2}M_{2}-x_{1}x_{2}M_{1}+N_{1}=0,\\ &x_{1}x_{2}M_{2}-x_{2}^{2}M_{1}+N_{2}=0.\end{split}$ (72)
Solutions of the system including equations (72) and (64) are presented in Item 1 of Table 1.
Consider now the tensor integrals of motion generated by $\mu_{(2)}^{ab}$ (39). Up to rotation transformations and arbitrary element $g({\bf x})$ we can restrict ourselves to the following version of nonzero parameters $\lambda^{ab}$:
@@
$M=\frac{G(\theta)+F(\varphi)}{\tilde{r}^{2}},\ N=F(\varphi)$
where $G(\theta)$ and $F(\varphi)$ are arbitrary functions of Euler angles, refer to Item 2 of Table 1 where the corresponding potential is represented also.
For some special form of the mentioned arbitrary functions represented in Item 2 of Table 1 the related PDM system admits more integrals of motion and is superintrgrable or even maximally superintegrable, see Items 1–3, of Table 2. One more superintegrable system which admit two integrals of motion of type (75) includes the arbitrary elements represented in the following formulae:
$\begin{split}&f=\frac{x_{1}^{2}x_{2}^{2}x_{3}^{2}r^{2}}{x_{3}^{2}(x_{3}^{2}(c_{1}x_{2}^{2}+c_{2}x_{1}^{2})+c_{1}x_{2}^{4}+c_{2}x_{1}^{4})+c_{4}\tilde{r}^{2}x_{1}^{2}x_{2}^{2}},\\ &V=\frac{x_{3}^{2}(x_{3}^{2}(c_{5}x_{2}^{2}+c_{6}x_{1}^{2})+c_{5}x_{2}^{4}+c_{6}x_{1}^{4})+c_{8}\tilde{r}^{2}x_{1}^{2}x_{2}^{2}}{x_{3}^{2}(x_{3}^{2}(c_{1}x_{2}^{2}+c_{2}x_{1}^{2})+c_{1}x_{2}^{4}+c_{2}x_{1}^{4})+c_{4}\tilde{r}^{2}x_{1}^{2}x_{2}^{2}},\end{split}$ (77)
while the related integrals of motion look as:
$\begin{split}&Q_{1}=L_{1}^{2}-\left(\frac{c_{2}x_{3}^{2}+c_{4}x_{2}^{2}}{x_{2}^{2}x_{3}^{2}}\cdot H\right)+\frac{c_{6}x_{3}^{2}+c_{8}x_{2}^{2}}{x_{2}^{2}x_{3}^{2}},\\ &Q_{2}=L_{2}^{2}-\left(\frac{c_{1}x_{3}^{4}+c_{2}x_{1}^{4}}{x_{1}^{2}x_{3}^{2}}\cdot H\right)+\frac{c_{5}x_{3}^{4}+c_{6}x_{1}^{4}}{x_{1}^{2}x_{3}^{2}}.\end{split}$ (78)
The case (74) is the most complicated and interesting. The related determining equations (52) is reduced to the following system:
@@
$\begin{split}&\omega=\sqrt{z^{2}+4\rho},\ \rho=\frac{\mu x_{1}^{2}-\nu x_{2}^{2}}{x_{3}^{2}}+\mu\nu(\mu-\nu),\\ &z=\frac{\nu x_{1}^{2}+\mu x_{2}^{2}+(\mu-\nu)(\mu^{2}-\nu^{2})x_{3}^{2}}{x_{3}^{2}\sqrt{\mu\nu(\mu-\nu))}}\end{split}$ (81)
@@
$M=\frac{F(\omega+z)+G(\omega-z)}{\omega x_{3}^{2}}.$ (82)
@@
Notice that linear combinations of solutions (84) also solve equations (79), (71). Just these results are presented in Item 5 of Table 1.
@@
Table 2. Inverse masses, potentials and integrals of motion for superintegrable scale invariant systems.
@@
1 $\frac{x_{3}^{2}\tilde{r}^{2}}{c_{1}\tilde{r}^{2}+F(\varphi)x_{3}^{2}}$ $\frac{G(\varphi)x_{3}^{2}+c_{2}\tilde{r}^{2}}{c_{1}\tilde{r}^{2}+F(\varphi)x_{3}^{2}}$ $\begin{array}[]{c}P_{3}^{2}+\frac{c_{2}}{x_{3}^{2}}-\left(\frac{c_{1}}{x_{3}^{2}}\cdot H\right),\\ K_{3}^{2}+\frac{c_{2}r^{4}}{x_{3}^{2}}-\left(\frac{c_{1}r^{4}}{x_{3}^{2}}\cdot H\right)+\hat{\eta},\\ L_{3}^{2}+G(\varphi)-(F(\varphi)\cdot H)\end{array}$
@@
3 $\frac{x_{3}^{2}x_{1}^{2}\tilde{r}}{c_{1}\tilde{r}x_{3}^{2}+c_{2}x_{2}x_{3}^{2}+c_{3}x_{1}^{2}\tilde{r}}$ $\frac{c_{4}\tilde{r}x_{3}^{2}+c_{5}x_{2}x_{3}^{2}+c_{6}x_{1}^{2}\tilde{r}}{c_{1}\tilde{r}x_{3}^{2}+c_{2}x_{2}x_{3}^{2}+c_{3}x_{1}^{2}\tilde{r}}$$\begin{array}[]{c}\{L_{3},P_{1}\}+\frac{2c_{4}x_{2}}{x_{1}^{2}}+\frac{c_{5}(\tilde{r}^{2}+x_{2}^{2})}{\tilde{r}x_{1}^{2}}\\ -\left(H\cdot(\frac{2c_{1}x_{2}}{x_{1}^{2}}+\frac{c_{2}(\tilde{r}^{2}+x_{2}^{2})}{\tilde{r}x_{1}^{2}})\right),\\ \{L_{3},K_{1}\}+\frac{2c_{4}x_{2}r^{2}}{x_{1}^{2}}+\frac{c_{5}r^{2}(\tilde{r}^{2}+x_{2}^{2})}{\tilde{r}x_{1}^{2}}\\ -\left(H\cdot(\frac{2r^{2}c_{1}x_{2}}{x_{1}^{2}}+\frac{c_{2}r^{2}(\tilde{r}^{2}+x_{2}^{2})}{\tilde{r}x_{1}^{2}})\right),\\ L_{3}^{2}-\left(\frac{c_{1}\tilde{r}^{2}+c_{2}x_{2}\tilde{r}}{x_{1}^{2}}\cdot H\right)+\frac{c_{4}\tilde{r}^{2}+c_{5}x_{2}\tilde{r}}{x_{1}^{2}},\\ P_{3}^{2}-\left(\frac{c_{3}}{x_{3}^{2}}\cdot H\right)+\frac{c_{6}}{x_{3}^{2}},\\ K_{3}^{2}-\left(\frac{c_{3}r^{4}}{x_{3}^{2}}\cdot H\right)+\frac{c_{6}r^{4}}{x_{3}^{2}}+\hat{\eta}\par\end{array}$
4 $\frac{x_{1}^{2}x_{2}^{2}x_{3}^{2}}{c_{1}x_{2}^{2}x_{3}^{2}+c_{2}x_{1}^{2}x_{3}^{2}+c_{3}x_{1}^{2}x_{2}^{2}}$ $\frac{c_{4}x_{2}^{2}x_{3}^{2}+c_{5}x_{1}^{2}x_{3}^{2}+c_{6}x_{1}^{2}x_{2}^{2}}{c_{1}x_{2}^{2}x_{3}^{2}+c_{2}x_{1}^{2}x_{3}^{2}+c_{3}x_{1}^{2}x_{2}^{2}}$ $\begin{array}[]{c}P_{2}^{2}-\left(\frac{c_{2}}{x_{2}^{2}}\cdot H\right)+\frac{c_{5}}{x_{2}^{2}},\\ P_{1}^{2}-\left(\frac{c_{1}}{x_{1}^{2}}\cdot H\right)+\frac{c_{4}}{x_{1}^{2}},\\ K_{2}^{2}-\left(\frac{c_{2}r^{4}}{x_{2}^{2}}\cdot H\right)+\frac{c_{5}r^{4}}{x_{2}^{2}}\\ +3r^{2}+15x_{2}^{2},\\ K_{1}^{2}-\left(\frac{c_{1}r^{4}}{x_{1}^{2}}\cdot H\right)+\frac{c_{4}r^{4}}{x_{1}^{2}},\\ +3r^{2}+15x_{1}^{2},\\ L_{1}^{2}-\left(\frac{c_{3}x_{2}^{4}+c_{2}x_{3}^{4}}{x_{2}^{2}x_{3}^{2}}\cdot H\right)+\frac{c_{6}x_{2}^{4}+c_{5}x_{3}^{4}}{x_{2}^{2}x_{3}^{2}},\\ L_{2}^{2}-\left(\frac{c_{3}x_{1}^{4}+c_{1}x_{3}^{4}}{x_{1}^{2}x_{3}^{2}}\cdot H\right)+\frac{c_{6}x_{1}^{4}+c_{3}x_{3}^{4}}{x_{1}^{2}x_{3}^{2}}\end{array}$
Thus we find all inequivalent PDM systems which are scale invariant and admit at least one second order integral of motion. They are represented in Tables 1 and 2 where $F(.),G(.),\tilde{F}(.),\tilde{G}(.)$ are arbitrary functions whose arguments are fixed in the brackets, $\omega,z,\rho$ and $\Phi_{m}$ are given by relations (81) and (83), $c_{1},c_{2},...$ are arbitrary real parameters. One more scale invariant and superintegrable PDM system is presented in equations (77) and (LABEL:equa) since the related arbitrary elements are too cumbersome to be placed in the table.
@@
Notice that systems whose arbitrary elements are present in Table 1 are integrable while the superintegrable systems are represented in Table 2 and equation ones are superintegrable. Moreover, Items 3 and 4 ot this table represent maximally superintegrable systems.
Notice that systems whose arbitrary elements are present in Table 1 are integrable while the superintegrable systems are represented in Table 2 and equation ones are superintegrable. Moreover, Items 3 and 4 ot this table represent maximally superintegrable systems.

### v3

@@ 3.5 Symmetries with differential terms $Q_{(n)}$ (22) quadratic in independent variables
The last class of symmetries which we are supposed to find for the scale invariant systems are integrals of motion (4) generated by Killing tensors (31). The specificity of this class is that it includes integrals of motions with different transformation properties with respect to the discrete symmetry (28). Namely, the vector integrals of motion generated by $\mu_{(1)}^{ab}$ are invariant with respect to transformations (28) while the tensor integrals of motion generated by $\mu_{(2)}^{ab}$ change their sign under this transformation. It means that in fact we have two independent subclasses of integrals of motion since the linear combinations of integrals of motion with different transformations properties with respect to symmetry transformations are forbidden.
3.5.1 Vector integrals of motion
@@
Let us start with vector integrals of motion which are generated by the Killing tensor $\mu_{(1)}^{ab}$ (31). Up to rotation transformations the only nonzero parameter in $\mu_{(1)}^{ab}$is $\lambda^{3}$, and the related equations (49) are reduced to the following form:
@@
In addition we have to take into account condition (35) while the analogue of equations (36) and (52) takes the following form:
$x_{a}N_{a}=c_{1}.$ (75)
Indeed, tensor $\mu_{(1)}^{ab}$ satisfies the condition $x_{c}\mu_{(1)c}^{ab}=2\mu_{(1)}^{ab}$, thus the related operator $Q_{(n)}$ (22) is not changed under the dilatation transformation. The same property can be requested to the total integral of motion (21) presented in (22). In other words, we can suppose that the first term $Q_{(0)}=P_{a}gP_{a}$ in (22) is not changed also. It is the case when function $g=\frac{N}{M}$ satisfies the same condition as $\mu_{(1)}^{ab}$, i.e., $x_{a}g_{a}=2g$. It is the case if $N$ satisfies condition (75) with $c_{1}=0$ since $M$ satisfies (35).
But what has happen in the case when parameter $c_{1}$ in (75) is not trivial? The related integral of motion will be changed under the conformal transformations, but this change is reduced to the form $Q\to Q+cH$ where $H$ is the Hamiltonian and $c$ is a constant. Such changes are surely admissible since $H$ commutes with itself.
The presented speculations are not necessary since equations (84) can be solved directly without using (75). However, the latter condition presents a nice tool to simplify calculations.
Multiplying the first of equations (84) by $x_{1}$, the second equation by $x_{2}$, the third by $x_{3}$, summing up the obtained results and using (75) we obtain the following consequence:
$r^{2}(x_{1}M_{2}-x_{2}M_{1})=-c_{1}.$ (76)
Its generic solution is
$M=-\frac{c_{1}\varphi+G(\theta)}{r^{2}}$ (77)
where $\varphi$ and $\theta$ are the Euler angles and $G(\theta)$ is an arbitrary function. This solution is valid for the initial equations (84) also.
3.5.2 Tensor integrals of motion
Consider now the tensor integrals of motion generated by $\mu_{(2)}^{ab}$ (31). Up to rotation transformations we can restrict ourselves to the following version of nonzero parameters $\lambda^{ab}$:
@@ 3.5.2 Tensor integrals of motion
$M=\frac{G(\theta)+F(\varphi)}{\tilde{r}^{2}},\ N=F(\varphi)$ (82)
where $G(\theta)$ and $F(\varphi)$ are arbitrary functions of Euler angles.
The case (79) is the most complicated and interesting. The related integral of motion (4) takes the form
$Q=\mu L_{1}^{2}+\nu L_{2}^{2}+Q_{(0)}+\eta$ (83)
and determining equations (49) are reduced to the following system:
@@
Without loss of generality parameters $\mu$ and $\nu$ are supposed to be non equal, i.e., $\mu\neq\nu$, since for $\mu=\nu$ operator (83) is proportional to a linear combination of the squared dilatation generator $D^{2}$ and operator (80) analysed in the above.
@@
$\begin{split}&\omega=\sqrt{z^{2}+4\rho},\ \rho=\frac{\mu x_{1}^{2}-\nu x_{2}^{2}}{x_{3}^{2}}+\mu\nu(\mu-\nu),\\ &z=\frac{\nu x_{1}^{2}+\mu x_{2}^{2}+(\mu-\nu)(\mu^{2}-\nu^{2})x_{3}^{2}}{x_{3}^{2}\sqrt{|\mu\nu(\mu-\nu)|}}\end{split}$ (86)
@@
$M=\frac{F(\omega+z)+G(\omega-z)}{x_{3}^{2}}.$ (87)
@@
Notice that linear combinations of solutions (82) also solve equations (84) and (75).
We find all inequivalent solutions of the determining equations (24). Thus we fix all inequivalent PDM systems which are scale invariant and admit at least one second order integral of motion. By definition such systems are integrable. However, some of them admit more than one integral of motion as it is shown in the following subsection.
3.5.3 Search for superintegrable sustems
We find all inequivalent solutions of the determining equations (24). Thus we fix all inequivalent PDM systems which are scale invariant and admit at least one second order integral of motion. By definition such systems are integrable.
The found mass functions $M$ include arbitrary coefficients and even arbitrary functions. For some special form of these arbitrary functions and special combinations of the values of arbitrary parameters the related PDM system can admit more integrals of motion and be superintrgrable or even maximally superintegrable. Just these possibilities are studied in the present section.
Let us consider step by step all solutions presented in formulae (40), (47), (62), (69), (72), (82), (87) and search for the cases when they satisfy the following functional equations:
$M=M^{\prime}$ (90)
where $M$ is the solution given by one of the mentioned formulae and $M^{\prime}$ is a solution given by another one.
Let us start with the case presented in (47). In this case we have three integrals of motion fixed in (48), two of which are functionally independent. One more integral of motion is the generator of conformal transformations $D$. Thus the related PDM system is superintegrable.
Comparing solutions (47) with (82) we observe that they coincide provided functions $F(\varphi)$ and $G(\theta)$ present in (82) have the following forms:
$F(\varphi)=\frac{c_{1}x_{2}^{2}}{x_{1}^{2}}+\frac{c_{2}x_{1}^{2}}{x_{2}^{2}}-c_{1}-c_{2},\ G(\theta)=\frac{c_{3}\tilde{r}^{2}}{x_{3}^{2}}.$ (91)
The related function $N$ is equal to $F(\varphi)$ given in (91) and integral of motion (80) is reduced to the following form:
$Q=L_{3}^{2}-\left(\frac{c_{1}x_{2}^{4}+c_{2}x_{1}^{4}}{x_{1}^{2}x_{2}^{2}}\cdot H\right)+\frac{c_{4}x_{2}^{4}+c_{5}x_{1}^{4}}{x_{1}^{2}x_{2}^{2}}.$ (92)
Since variables $x_{1},x_{2}$ and $x_{3}$ are involved into formulae (47) in completely symmetric way we can conclude that there exist two more integrals of motion admitted by the considered system which can be obtained from (92) by the changes $L_{3}\to L_{1},x_{3}\to x_{1},x_{2}\to x_{2},x_{1}\to-x_{3}$ and $L_{3}\to L_{2},x_{1}\to x_{1},x_{2}\to x_{3},x_{3}\to-x_{2}$.
Let us compare solutions for $M$ given by relations (40), (62), (72) and (82). We recognize that functions (40) and (62) are particular cases of the function presented in (82) with $G(\theta)=\tan(\theta)^{2}$ and $G(\theta)=\cot(\theta)$ respectively. The same is true for solution (72) which has the form (82) (and (40)!) with $G(\theta)=c_{3}\tan(\theta)^{2}$ and $F(\varphi)=\frac{c_{1}}{\cos(\varphi)^{2}}+\frac{c_{2}\tan(\varphi)}{\cos(\theta)}$. On the other hand functions (40) and (62) are essentially different and cannot be reduced one to another for some particular form of arbitrary functions or nontrivial arbitrary constants. The case of the trivial constants $c_{1}$ is forbidden since in this case we have as minimum two parametric Lie symmetry group.
It follows from the above that the PDM systems whose masses have the form given by equations (62), (72) and (82) admit the additional integral of motion presented in (42) and so are superintegrable. Moreover, the mass function (72) is maximally superintegrable since in addition it admits one more integral of motion whose form is presented in (80).
The mass functions presented in (69) and (85) have rather specific dependence on the angular variables which has nothing to do with the other found versions of such functions. Any of the related PDN systems admits only one integral of motion additional to $D$ and is integrable but not superintegrable.
The only version which was not considered jet is given by equation (87). It is extremely difficult to look directly for the cases when it can coincide with one of the version presented in (40), (47), (62), (69), (72), (82), (87). However we can search for the cases when equations (85) are compatible with equations (51) corresponding to these versions. As a result we find the only case when it is possible, and this case corresponds to the following particular form of functions (87) and (82) which coincide:
$M=\frac{c_{1}}{x_{1}^{2}}+\frac{c_{2}}{x_{2}^{2}}+\frac{c_{3}}{x_{3}^{2}}+\frac{c_{4}}{r^{2}}$
while the related potential (25) looks as follows:
$V=\frac{r^{2}(c_{5}x_{2}^{2}x_{3}^{2}+c_{6}x_{1}^{2}x_{3}^{2}+c_{7}x_{1}^{2}x_{2}^{2})+c_{8}x_{1}^{2}x_{2}^{2}x_{3}^{2}}{r^{2}(c_{1}x_{2}^{2}x_{3}^{2}+c_{2}x_{1}^{2}x_{3}^{2}+c_{3}x_{1}^{2}x_{2}^{2})+c_{4}x_{1}^{2}x_{2}^{2}x_{3}^{2}}$
The corresponding integrals of motion are given by equation (80) and where $N=\frac{c_{1}}{\cos(\varphi)^{2}}+\frac{c_{2}}{\sin(\varphi)^{2}}$ and $\tilde{N}=\frac{c_{5}}{\cos(\varphi)^{2}}+\frac{c_{6}}{\sin(\varphi)^{2}}.$
3.5.4 Classification results
Thus we find all inequivalent position dependent masses which correspond to systems which are scale invariant and admit at least one second order integral of motion.
In this section we summarize the obtained results. They are represented in Table 1 where $F(.),G(.),\tilde{F}(.),\tilde{G}(.)$ are arbitrary functions whose arguments are fixed in the brackets, $\omega,z,\rho$ and $\Phi_{m}$ are given by relations (86) and (88), $\hat{\eta}^{a}=3r^{2}+15x_{a}^{2},a=1,2,3$, $W=x_{1}^{2}x_{2}^{2}x_{3}^{2},$ $c_{1},c_{2},...$ are arbitrary real parameters. In addition, notation (43) is used.
@@ 3.5.4 Classification results
Table 1. Inverse masses, potentials and integrals of motion for scale invariant systems.
@@
3 $\frac{\tilde{r}^{4}}{c_{1}(x_{3}^{2}+r^{2})e^{\pm 2\varphi}+c_{2}x_{3}\tilde{r}e^{\pm\varphi}+c_{3}\tilde{r}^{2}}$ $\frac{c_{4}(x_{3}^{2}+r^{2})e^{\pm 2\varphi}+c_{5}x_{3}\tilde{r}e^{\pm\varphi}+c_{6}\tilde{r}^{2}}{c_{1}(x_{3}^{2}+r^{2})e^{\pm 2\varphi}+c_{2}x_{3}\tilde{r}e^{\pm\varphi}+c_{3}\tilde{r}^{2}}$$\begin{array}[]{c}\{P_{3},(D\mp L_{3})\}\\ +2\left((\frac{c_{2}e^{\pm\varphi}}{\tilde{r}}+\frac{c_{1}x_{3}e^{\pm 2\varphi}}{\tilde{r}^{2}})\cdot H\right)\\ -2(\frac{c_{5}e^{\pm\varphi}}{\tilde{r}}+\frac{c_{4}x_{3}e^{\pm 2\varphi}}{\tilde{r}^{2}}),\\ \{K_{3},(D\mp L_{3})\}-15x_{3}\\ +2\left(r^{2}(\frac{c_{2}e^{\pm\varphi}}{\tilde{r}}+\frac{c_{1}x_{3}e^{\pm 2\varphi}}{\tilde{r}^{2}})\cdot H\right)\\ -2r^{2}(\frac{c_{5}e^{\pm\varphi}}{\tilde{r}}+\frac{4_{6}x_{3}e^{\pm 2\varphi}}{\tilde{r}^{2}})\end{array}$
6 $\frac{x_{3}^{2}\tilde{r}^{2}}{c_{1}\tilde{r}^{2}+F(\varphi)x_{3}^{2}}$ $\frac{G(\varphi)x_{3}^{2}+c_{2}\tilde{r}^{2}}{c_{1}\tilde{r}^{2}+F(\varphi)x_{3}^{2}}$ $\begin{array}[]{c}P_{3}^{2}+\frac{c_{2}}{x_{3}^{2}}-\left(\frac{c_{1}}{x_{3}^{2}}\cdot H\right),\\ K_{3}^{2}+\frac{c_{2}r^{4}}{x_{3}^{2}}-\left(\frac{c_{1}r^{4}}{x_{3}^{2}}\cdot H\right)+\hat{\eta}^{3},\\ L_{3}^{2}+G(\varphi)-(F(\varphi)\cdot H)\end{array}$
@@
8 $\frac{r^{2}W}{r^{2}(c_{1}x_{2}^{2}x_{3}^{2}+c_{2}x_{1}^{2}x_{3}^{2}+c_{3}x_{1}^{2}x_{2}^{2})+c_{4}W}$ $\frac{r^{2}(c_{5}x_{2}^{2}x_{3}^{2}+c_{6}x_{1}^{2}x_{3}^{2}+c_{7}x_{1}^{2}x_{2}^{2})+c_{8}W}{r^{2}(c_{1}x_{2}^{2}x_{3}^{2}+c_{2}x_{1}^{2}x_{3}^{2}+c_{3}x_{1}^{2}x_{2}^{2})+c_{4}W}$ $\begin{array}[]{c}L_{1}^{2}-\left(\frac{c_{3}x_{2}^{4}+c_{2}x_{3}^{4}}{x_{2}^{2}x_{3}^{2}}\cdot H\right)+\frac{c_{6}x_{2}^{4}+c_{5}x_{3}^{4}}{x_{2}^{2}x_{3}^{2}},\\ L_{2}^{2}-\left(\frac{c_{3}x_{1}^{4}+c_{1}x_{3}^{4}}{x_{1}^{2}x_{3}^{2}}\cdot H\right)+\frac{c_{6}x_{1}^{4}+c_{3}x_{3}^{4}}{x_{1}^{2}x_{3}^{2}}\end{array}$
9 $\frac{x_{3}^{2}x_{1}^{2}\tilde{r}}{c_{1}\tilde{r}x_{3}^{2}+c_{2}x_{2}x_{3}^{2}+c_{3}x_{1}^{2}\tilde{r}}$ $\frac{c_{4}\tilde{r}x_{3}^{2}+c_{5}x_{2}x_{3}^{2}+c_{6}x_{1}^{2}\tilde{r}}{c_{1}\tilde{r}x_{3}^{2}+c_{2}x_{2}x_{3}^{2}+c_{3}x_{1}^{2}\tilde{r}}$$\begin{array}[]{c}\{L_{3},P_{1}\}+\frac{2c_{4}x_{2}}{x_{1}^{2}}+\frac{c_{5}(\tilde{r}^{2}+x_{2}^{2})}{\tilde{r}x_{1}^{2}}\\ -\left(H\cdot(\frac{2c_{1}x_{2}}{x_{1}^{2}}+\frac{c_{2}(\tilde{r}^{2}+x_{2}^{2})}{\tilde{r}x_{1}^{2}})\right),\\ \{L_{3},K_{1}\}+\frac{2c_{4}x_{2}r^{2}}{x_{1}^{2}}+\frac{c_{5}r^{2}(\tilde{r}^{2}+x_{2}^{2})}{\tilde{r}x_{1}^{2}}\\ -\left(H\cdot(\frac{2r^{2}c_{1}x_{2}}{x_{1}^{2}}+\frac{c_{2}r^{2}(\tilde{r}^{2}+x_{2}^{2})}{\tilde{r}x_{1}^{2}})\right),\\ L_{3}^{2}-\left(\frac{c_{1}\tilde{r}^{2}+c_{2}x_{2}\tilde{r}}{x_{1}^{2}}\cdot H\right)+\frac{c_{4}\tilde{r}^{2}+c_{5}x_{2}\tilde{r}}{x_{1}^{2}},\\ P_{3}^{2}-\left(\frac{c_{3}}{x_{3}^{2}}\cdot H\right)+\frac{c_{6}}{x_{3}^{2}},\\ K_{3}^{2}-\left(\frac{c_{3}r^{4}}{x_{3}^{2}}\cdot H\right)+\frac{c_{6}r^{4}}{x_{3}^{2}}+\hat{\eta}^{3}\par\end{array}$
10 $\frac{x_{1}^{2}x_{2}^{2}x_{3}^{2}}{c_{1}x_{2}^{2}x_{3}^{2}+c_{2}x_{1}^{2}x_{3}^{2}+c_{3}x_{1}^{2}x_{2}^{2}}$ $\frac{c_{4}x_{2}^{2}x_{3}^{2}+c_{5}x_{1}^{2}x_{3}^{2}+c_{6}x_{1}^{2}x_{2}^{2}}{c_{1}x_{2}^{2}x_{3}^{2}+c_{2}x_{1}^{2}x_{3}^{2}+c_{3}x_{1}^{2}x_{2}^{2}}$ $\begin{array}[]{c}P_{2}^{2}-\left(\frac{c_{2}}{x_{2}^{2}}\cdot H\right)+\frac{c_{5}}{x_{2}^{2}},\\ P_{1}^{2}-\left(\frac{c_{1}}{x_{1}^{2}}\cdot H\right)+\frac{c_{4}}{x_{1}^{2}},\\ K_{2}^{2}-\left(\frac{c_{2}r^{4}}{x_{2}^{2}}\cdot H\right)+\frac{c_{5}r^{4}}{x_{2}^{2}}+\hat{\eta}^{2},\\ K_{1}^{2}-\left(\frac{c_{1}r^{4}}{x_{1}^{2}}\cdot H\right)+\frac{c_{4}r^{4}}{x_{1}^{2}}+\hat{\eta}^{1},\\ L_{1}^{2}-\left(\frac{c_{3}x_{2}^{4}+c_{2}x_{3}^{4}}{x_{2}^{2}x_{3}^{2}}\cdot H\right)+\frac{c_{6}x_{2}^{4}+c_{5}x_{3}^{4}}{x_{2}^{2}x_{3}^{2}},\\ L_{2}^{2}-\left(\frac{c_{3}x_{1}^{4}+c_{1}x_{3}^{4}}{x_{1}^{2}x_{3}^{2}}\cdot H\right)+\frac{c_{6}x_{1}^{4}+c_{3}x_{3}^{4}}{x_{1}^{2}x_{3}^{2}}\end{array}$
@@
Notice that the systems whose arbitrary elements are present items 1-5 are integrable while while the remaining items represent superintegrable systems. Moreover, the systems fixed in Items 9 and 10 are maximally superintegrable.
A bit mysterious system presented in Item 10 admits as much as seven linearly independent integrals of motion, namely,
$\begin{split}&Q_{1}=P_{1}^{2}-\left(\frac{c_{1}}{x_{1}^{2}}\cdot H\right)+\frac{c_{4}}{x_{1}^{2}},\\ &Q_{2}=P_{2}^{2}-\left(\frac{c_{2}}{x_{2}^{2}}\cdot H\right)+\frac{c_{5}}{x_{2}^{2}},\\ &Q_{3}=L_{1}^{2}-\left(\frac{c_{3}x_{2}^{4}+c_{2}x_{3}^{4}}{x_{2}^{2}x_{3}^{2}}\cdot H\right)+\frac{c_{6}x_{2}^{4}+c_{5}x_{3}^{4}}{x_{2}^{2}x_{3}^{2}},\\ &Q_{4}=L_{2}^{2}-\left(\frac{c_{3}x_{1}^{4}+c_{1}x_{3}^{4}}{x_{1}^{2}x_{3}^{2}}\cdot H\right)+\frac{c_{6}x_{1}^{4}+c_{3}x_{3}^{4}}{x_{1}^{2}x_{3}^{2}},\\ &Q_{5}=K_{1}^{2}-\left(\frac{c_{1}r^{4}}{x_{1}^{2}}\cdot H\right)+\frac{c_{4}r^{4}}{x_{1}^{2}}+\hat{\eta}^{1},\\ &Q_{6}=K_{2}^{2}-\left(\frac{c_{2}r^{4}}{x_{2}^{2}}\cdot H\right)+\frac{c_{5}r^{4}}{x_{2}^{2}}+\hat{\eta}^{2},\\ &Q_{7}=D.\end{split}$ (93)
Operators $Q_{5}$ and $Q_{6}$ are connected with $Q_{1}$ and $Q_{2}$ via discrete transformation (28). Since the maximal number of integrals of motion allowed for Hamiltonian systems with three degrees of freedom is equal to 4, the remaining integrals of motion $Q_{1}-Q_{4}$ and $Q_{5}$ should be functionally dependent. This dependence is implicit and can be found by calculation the double commutator $[Q_{1}-Q_{2},[Q_{1}-Q_{2},Q_{3}+Q_{4}]]$.